OpenStudy (anonymous):

if a polynomial function f(x) has roots 4-13i and 5, what must be a factor of f(x) a. (x+(13-4i)) b. (x-(13+4i)) c. (x+(4+13i)) d. (x-(4+13i))

OpenStudy (solomonzelman):

when your roots are "a" and "-b" the factors of the function are, (x-a) and (x+b) and function is as follows, f(x) = (x-a)(x+b)

OpenStudy (solomonzelman):

"x minus the additive inverse of the root" this is inside the parenthesis.

OpenStudy (solomonzelman):

do you want to see an example similar to yours ?

OpenStudy (anonymous):

no thanks I got it thank you

OpenStudy (solomonzelman):

I think that your answers are not correct. It is supposed to be a product (multiplication) of factors, so that a root would make one part of the product a zero, so that you get a zero for f(x) (when you plug in the value of the root instead of x) Your choices however are either wrong, or have some typo....

OpenStudy (solomonzelman):

For roots "d" and "f" it is supposed to be \(\large\color{black}{f(x)=(x-d)(x-f) }\) in other words, \(\large\color{black}{f(x)=(x-d) \color{blue}{\LARGE\times}(x-f) }\) Not, \(\large\color{black}{f(x)=(x-d) \color{red}{-}(x-f) }\) and NOT \(\large\color{black}{f(x)=(x-d) \color{red}{+}(x-f) }\) and not even \(\large\color{black}{f(x)=(x-d) \color{red}{\div}(x-f) }\) (in this scenario, the "f" is when function is undefined, not a root.

OpenStudy (solomonzelman):